A Bibliography of Publications of William M. Coughran

William M. Coughran AT&T Bell Laboratories

600 Mountain Av., Rm. 2T-502

Murray Hill, NJ 07974-0636 USA

Tel: +1 908 582 6619 FAX: ?n/a?

E-mail: wmc@research.bell-labs.com (Internet)

13 October 2017 Version 1.21

Abstract

This bibliography records publications of William M. Coughran.

Title word cross-reference

C_{2}H_{4} [CRSM73]. IV [CFS94]. N_{2} [CRSM73].

/ [Boi97].

Adaptive [BCF91, CS90, CPS91]. Advanced [BDR^{+}94]. Algorithm [CH81]. Algorithms [BBC^{+}90, CJ90, CGP^{+}92b, CFS94]. Alternate [BCCS89]. Alternate-Block-Factorization [BCCS89]. Analysis [BCC^{+}79, Cou77, CGR87, CJ90]. Analyzer [CGR83]. Animation [CG90c]. Applied [BCG94a, BCG94b]. Approach [LFC90]. Approximate [Cou80]. Aspects [BCF^{+}86, CGR87, SCF^{+}86]. assessment [Boi97]. Augmented [CFS94].

based [BGP^{+}92, Cou97]. Bénard [BCG^{+}84, GCS82, GC84]. BiCMOS [PBR^{+}92]. Bifurcations [CGP92a]. Bipolar [PBR^{+}92]. Block [BCCS89]. Boundary [Cou80, Cou84, Cou85].

CAzM [CGR83, CEK^{+}89]. Cell [GCC88]. Challenges [SC94]. Chaos [GCC88]. Characteristics [CGP92a]. Characterization [PBR^{+}92]. Charges [CFG89]. Circuit [BDR^{+}94, CGR83, CGR87, CFS94, LFC90]. Circuit-Analyzer [CGR83]. Circuits [BCF^{+}85]. CMOS [CPS88]. CO [CRSM73]. Collaboration [CEK^{+}89]. Combined [BDR^{+}94]. Complete [CH81]. computation [Cou97]. Computational [BCF^{+}86, CGR87, CPS88, CGP^{+}92b, SC94]. Computing [CG89, CGP92a]. Concerning [Cou77]. Conditions [Cou84, Cou85].

Conference [Boi97]. Construct [CH81]. Construction [Cou77]. Continuation [CPS89, CS90, CFS94]. Convecting [CG87]. Convection [GCS82, GCC88, GC84]. Convective [BPC^{+}91]. Coupled [BGP^{+}92, BCG94a, BCG94b]. Current [CGP92a]. Current-Voltage [CGP92a]. Curves [CRSM73].

Decomposition [BCG94a, BCG94b]. Design [CEK^{+}89, CFS86]. Development [CFS86]. Device [BCF^{+}86, BCD^{+}89, BBC^{+}90, BDR^{+}94, CFG89, CS90, CJ90, CPS91, CFS94, CFS86, PCR^{+}91, PBR^{+}92, SCF^{+}86, CPS89]. Device-Circuit [BDR^{+}94]. Devices [BCF^{+}85, BCPS92, BDR^{+}94, CGP92a, SC94]. Differential [BCCS89]. Diffusion [BCF91]. dimensional [PBR^{+}92]. diminishing [CGR86]. Discrete [Cou84, Cou85]. Display [CG92]. Domain [BCG94a, BCG94b]. Drift [BCF91]. Drift-Diffusion [BCF91]. Dynamic [CG91].

Editor [CG83]. Effects [BCPS92]. Electrical [LFC90]. Element [CFS94]. Energy [BPC^{+}91, CRSM73]. enhancement [Boi97]. Equation [BCG^{+}84]. Equations [BCCS89, BCG94a, BCG94b, BCF91, CRSM73, CFS94]. Equations-of-motion [CRSM73]. Evaluation [BPC^{+}91]. Extracting [CFG89].

Factorization [BCCS89]. Faster [CS90]. Fitting [CFG89]. Flows [CG87, GCC88]. Folds [CGP92a]. Formation [CG87, GCS82, GC84]. Fully [BPC^{+}91, BGP^{+}92]. Functions [CG92].

Generation [CPS91]. Gradient [CFG89]. Grid [BCF91, CPS91]. Grove [CG83]. Guide [BCC^{+}79].

HBTs [BPC^{+}91, BGP^{+}92]. Hearing [CG91]. Hierarchical [BCPS92]. Hydrodynamic [BPC^{+}91, BGP^{+}92]. Hyperbolic [Cou80, Cou84, Cou85].

IFIP [Boi97]. II [CCLW94]. Incorporating [CFS94]. Inferno [Cou97]. Influence [BPC^{+}91]. Initial [Cou80, Cou84, Cou85]. Initial-Boundary [Cou80]. Initial-Boundary-Value [Cou84, Cou85]. InP [BGP^{+}92]. InP-based [BGP^{+}92]. Instability [BCG^{+}84]. Integrated [PBR^{+}92]. Interpolating [CH81]. Iteration [CJ90]. Iterative [BCD^{+}89].

July [Boi97].

Kingdom [Boi97].

Language [GC83]. Languages [CG79]. Large [GCC88]. Large-Cell [GCC88]. Latchup [CPS88]. Library [BCC^{+}79, CCGH80, Cou77]. Loess [CG91].

Macromodeling [CGR83]. Mean [GCC88]. Meshes [CS90]. Method [CRSM73]. Methods [BCD^{+}89, CPS88, CPS89]. Microelectronics [CGP^{+}92b]. Mixed [LFC90]. Mixed-Mode [LFC90]. Mode [LFC90]. Model [BPC^{+}91, BCG^{+}84]. Modeling [CS90, CG90b]. Models [CFS94]. Modular [CJ90]. motion [CRSM73].

NAPLUG [BCC^{+}79]. Near

[BCG^{+}84, GCS82, GC84]. Network [Cou97]. Network-based [Cou97]. Noncharacteristic [Cou84, Cou85]. Nonequilibrium [BCPS92]. Nonlinear [GCS82, GC84]. Note [Cou77]. Numerical [BCG^{+}84, BCC^{+}79, CCGH80, Cou77, CG79, CG87, Boi97].

Object [LFC90]. Object-Oriented

[LFC90]. Onset [BCG^{+}84, GCS82, GCC88, GC84]. Oriented [LFC90]. Outer [CJ90]. Oxford [Boi97].

Parallel [BCG94a, BCG94b]. Part [CJ90]. Partial [BCCS89]. Parts [CCLW94]. Pattern [CG87, GCS82, GC84]. PDE [BCPS92]. Industry [CEK^{+}89]. Philosophy [CG89]. Pine [GC83]. Polygons [CG92]. Potential [CRSM73]. Problems [Cou80, Cou84, Cou85]. Procedure [BCCS89]. proceedings [Boi97]. Process [PBR^{+}92]. Program [BCC^{+}79, Cou77]. Programming [GC83]. Progress [BBC^{+}90].

Quality [Boi97].

Rayleigh [BCG^{+}84, GCS82, GC84]. Rayleigh-Bénard [BCG^{+}84, GCS82, GC84]. Refinement [BCF91].

Scientific [CG89, CG90a, CG90c, Cou97]. Seeing [CG91]. Semiconductor [BCD^{+}89, BBC^{+}90, BCPS92, BDR^{+}94, CJ90, CGP92a, SCF^{+}86, SC94, CPS89]. Semiconductors [CCLW94]. Shaded [CG92]. Sight [CG90a]. Silicon [BCF^{+}85, CG90b, PCR^{+}91]. Simulation [BCF^{+}85, BCF^{+}86, BCD^{+}89, BBC^{+}90, BGP^{+}92, BCPS92, BDR^{+}94, CPS88, CJ90, CPS91, CFS94, LFC90, PCR^{+}91, PBR^{+}92, SCF^{+}86, CGR86, CPS89]. Simulations [CFG89, SC94]. Software [CG79, CEK^{+}89, Boi97]. Solution [BCG^{+}84, Cou80]. Some [SCF^{+}86]. Sound [CG90a, CG92]. Space [CG92]. Spatial [CS90]. Splines [CH81, CGR86]. State [CPS88]. Steady [CPS88]. Steady-State [CPS88]. Strategy [BCF91]. Studies [CG87, CG90b]. Submicron [PBR^{+}92]. Supercomputers [CFS86]. Support [CCGH80]. Surfaces [CG91]. Systems [BCCS89].

TC2 [Boi97]. TC2/WG2.5 [Boi97]. Techniques [CG90c]. Technology [CFS86, PBR^{+}92]. Thermal [BGP^{+}92]. Three [CG92, PBR^{+}92]. Three-dimensional [PBR^{+}92]. Time [CG92]. Tools [CG89]. Transient [BCF^{+}85, BCF^{+}86, CJ90]. Transistor [CFG89]. Transistors [PBR^{+}92]. Transport [BCPS92, BCG94a, BCG94b].

ULSI [PCR^{+}91, SC94]. United [Boi97]. University [CEK^{+}89]. University/ Industry [CEK^{+}89]. User [BCC^{+}79]. Using [BPC^{+}91, CG92, PBR^{+}92, CS90].

Value [Cou80, Cou84, Cou85]. Variables [CG92]. Variation [CGR86]. via [Cou97]. Video [CG90a]. VLSI [CPS91, CFS86]. Voltage [CGP92a].

WG2.5 [Boi97]. Working [Boi97].

References

Bank:1990:RPA

[BBC^{+}90] R. E. Bank, J. Bürgler, W. M. Coughran, Jr., W. Fichtner, and R. K. Smith. Recent progress in algorithms for semiconductor device simulation. In R. E. Bank, R. Bulirsch, and K. Merten, editors, Mathematical Modelling and Simulation of Electrical Circuits and Semiconductor Devices, pages 125–140. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1990.

Bolstad:1979:NAP

[BCC^{+}79] J. H. Bolstad, T. F. Chan, W. M. Coughran, Jr., W. D. Gropp, E. H. Grosse, M. T. Heath, R. J. LeVeque, F. T. Luk, S. G. Nash, and L. N. Trefethen. Numerical analysis program library user’s guide (NAPLUG). User Note 82, SLAC Computing Services, 1979. First issued in 1976 by Chan, Coughran, Heath, and Luk.

Bank:1989:ABF

[BCCS89] R. E. Bank, T. F. Chan, W. M.

Coughran, Jr., and R. K. Smith. The alternate-block-factorization procedure for systems of partial differential equations. BIT, 29(4): 938–954, December 1989. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Bank:1989:IMS

[BCD^{+}89] R. E. Bank, W. M. Coughran, Jr., M. A. Driscoll, R. K. Smith, and W. Fichtner. Iterative methods in semiconductor device simulation. Comput. Phys. Comm, 53 (1–3):201–212, May 1989. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0010465589901604.

Bank:1985:TSS

[BCF^{+}85] R. E. Bank, W. M. Coughran, Jr., W. Fichtner, E. H. Grosse, D. J. Rose, and R. K. Smith. Transient simulation of silicon devices and circuits. IEEE Transactions on Computer-Aided Design, CAD-4: 436–451, 1985. Simultaneously published in IEEE Transactions on Electron Devices ED-32.

Bank:1986:CATa

[BCF^{+}86] R. E. Bank, W. M. Coughran, Jr., W. Fichtner, D. J. Rose, and R. K. Smith. Computational aspects of transient device simulation. In W. L. Engl, editor, Process and Device Simulation, pages 229–264. North Holland-Elsevier Science Publishers, Amsterdam, New York, Oxford, 1986.

Buergler:1991:AGRa

[BCF91] J. Bürgler, W. M. Coughran, Jr., and W. Fichtner. An adaptive grid refinement strategy for the drift-diffusion equations. IEEE Transactions on Computer-Aided Design, 10:1251–1258, 1991.

Bjorstad:1984:NSMa

[BCG^{+}84] P. E. Bjørstad, W. M. Coughran, Jr., H. S. Greenside, D. J. Rose, and N. L. Schryer. Numerical solution of a model equation near the onset of the Rayleigh-Bénard instability. In G. Birkhoff and A. Schoenstadt, editors, Elliptic Problem Solvers II, pages 531– 543. Academic Press, New York, NY, USA, 1984.

Bjorstad:1994:PDDa

[BCG94a] P. E. Bjørstad, W. M. Coughran, Jr., and E. H. Grosse. Parallel domain decomposition applied to coupled transport equations. Numerical Analysis Manuscript 94-3, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1994.

                                        URL ftp://netlib.bell-labs.
                                        com/netlib/att/cs/doc/94/4-
                                        03.ps.gz.

Bjorstad:1994:PDDb

[BCG94b] Petter Bjørstad, W. M. Coughran, Jr., and Eric Grosse. Parallel domain decomposition applied to coupled transport equations. In David E. Keys and Jinchao Xu, editors, Domain Decomposition Methods in Scientific and Engineering Computing, pages 369–380. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-5171-3. LCCN QA402.2 I55 1993.

URL ftp://cm.bell-labs.com/
cm/cs/doc/94/4-03.ps.gz.

Benvenuti:1992:HPSa

[BCPS92] A. Benvenuti, W. M. Coughran, Jr., M. R. Pinto, and N. L. Schryer. Hierarchical PDE simulation of nonequilibrium transport effects in semiconductor devices. In Workshop on Numerical Modeling of Processes and Devices for Integrated Circuits: NUPAD IV. Technical Digest, pages 155–160. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. Extended abstract.

Buergler:1994:CDCa

[BDR^{+}94] J. F. Bürgler, H. Dettmer, C. Riccobene, W. M. Coughran, Jr., and W. Fichtner. Combined device-circuit simulation for advanced semiconductor devices. In W. M. Coughran, Jr., J. Cole, P. Lloyd, and J. White, editors, Semiconductors, Part II, pages 89–107. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1994. ISBN 0-387-94251-3. Proceedings of the IMA 1991 Summer Program on Semiconductors, Minneapolis.

Benvenuti:1992:CTFa

[BGP^{+}92] A. Benvenuti, G. Ghione, M. R. Pinto, W. M. Coughran, Jr., and N. L. Schryer. Coupled thermal, fully hydrodynamic simulation of InP-based HBTs. In IEDM Technical Digest ’92, pages 737–740. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992.

Boisvert:1997:QNS

[Boi97] R. F. Boisvert, editor. Quality of numerical software: assessment and enhancement / proceedings of the IFIP TC2/WG2.5 Working Conference on the Quality of Numerical Software, Assessment and Enhancement, Oxford, United Kingdom, 8–12 July 1996. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

Benvenuti:1991:EICa

[BPC^{+}91] A. Benvenuti, M. R. Pinto, W. M. Coughran, Jr., N. L. Schryer, C. U. Naldi, and G. Ghione. Evaluation of the influence of convective energy in HBTs using a fully hydrodynamic model. In IEDM Technical Digest ’91, pages 499– 502, 1991.

Chan:1980:NLS

[CCGH80] Tony F. Chan, William M. Coughran, Jr., Eric H. Grosse, and Michael T. Heath. A numerical library and its support.

ACM Transactions on Mathematical Software, 6(2):135–145, June 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coughran:1994:SIII

[CCLW94] W. M. Coughran, Jr., J. Cole, P. Lloyd, and J. White, editors. Semiconductors, Parts I and II. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1994. ISBN 0-387-94250-5 and 0-387-94251-3. Proceedings of the IMA 1991 Summer Program on Semiconductors, Minneapolis.

Coughran:1989:CUIC

[CEK^{+}89] W. M. Coughran, Jr., D. J. Erdman, S. W. Kenkel, G. B. Nifong, D. J. Rose, and R. Subrahmanyan. CAzM: University/industry collaboration in design software. In Eighth Biennial University-Government-Industry Microelectronics Symposium, pages 102–105, 1989.

Coughran:1989:ETC

[CFG89] W. M. Coughran, Jr., W. Fichtner, and E. H. Grosse. Extracting transistor charges from device simulations by gradient fitting. IEEE Transactions on Computer-Aided Design, 8:380–394, 1989.

Hess:1986:SVT

[CFS86] W. M. Coughran, Jr., W. Fichtner, and R. K. Smith. Supercomputers in VLSI technology development and device design. In K. Hess, editor, Large Scale Computational Device Modeling, page 175. University of Illinois Press, 1986. Abstract.

Coughran:1994:AADSM

[CFS94] W. M. Coughran, Jr., R. W. Freund, and R. K. Smith. Algorithms for augmented device simulation models incorporating IV continuation and circuit element equations. In Workshop on Numerical Modeling of Processes and Devices for Integrated Circuits: NUPAD V. Technical Digest, pages 127–128. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. Extended abstract.

Coughran:1979:NLNS

[CG79] W. M. Coughran, Jr. and E. H. Grosse. New languages for numerical software. SIGNUM Newsletter, 14:73–75, 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Coughran:1983:GE

[CG83] W. M. Coughran, Jr. and E. H. Grosse. The grove editor. Numerical Analysis Manuscript 83-3, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1983.

Coughran:1987:PFCF

[CG87] W. M. Coughran, Jr. and H. S. Greenside. Numerical studies of pattern formation in convecting flows. SIAM News 20, pages 8– 9, March 1987.

Coughran:1989:PSC

[CG89] W. M. Coughran, Jr. and Eric Grosse. A philosophy for scientific computing tools. SIGNUM Newsletter, 24(2/3): 2–9, April/July 1989. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Coughran:1990:SVSS

[CG90a] W. M. Coughran, Jr. and E. H.

Grosse. Scientific video: Sight and sound. Scientific Video 90-2, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1990. Videotape.

Coughran:1990:SMS

[CG90b] W. M. Coughran, Jr. and E. H.

Grosse. Silicon modeling studies. Scientific Video 90-1, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1990. Videotape.

Farrell:1990:TSA

[CG90c] W. M. Coughran, Jr. and Eric Grosse. Techniques for scientific animation. In E. Farrell, editor, Proceedings of the SPIE, volume 1259, pages 72–79. SPIE Optical Engineering Press, Bellingham, WA, USA, 1990. Associated videotape in 1259-V collection.

Coughran:1991:SHD

[CG91] W. M. Coughran, Jr. and Eric Grosse. Seeing and hearing dynamic loess surfaces. In Interface’91 Proceedings, pages 224– 228. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1991.

URL ftp://netlib.bell-labs.
com/netlib/att/cs/doc/91/4-
07.ps.gz.

Gaffney:1992:DFT

[CG92] W. M. Coughran, Jr. and Eric Grosse. Display of functions of three space variables and time using shaded polygons and sound. In P. W. Gaffney and E. N. Houstis, editors, Programming Environments for High-Level Scientific Problem Solving, pages 271–276. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1992. ISBN 0-444-89176-

5. URL ftp://netlib.bell-
labs.com/netlib/att/cs/doc/
91/4-09.ps.gz. IFIP TC 2 WG

2.5 Proceedings of Working Conference on Programming Environments for High-Level Scientific Problem Solving, Karlsruhe.

Coughran:1992:CFB

[CGP92a] W. M. Coughran, Jr., E. H. Grosse, and M. R. Pinto. Computing folds and bifurcations in current-voltage characteristics of semiconductor devices. In Workshop on Numerical Modeling of Processes and Devices for Integrated Circuits: NUPAD IV. Technical Digest, pages 149–153. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. Extended abstract.

Coughran:1992:ACM

[CGP^{+}92b] W. M. Coughran, Jr., E. H. Grosse, M. R. Pinto, C. S. Rafferty, N. L. Schryer, and R. K. Smith. Algorithms for computational microelectronics. Scientific Video 92-1, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1992. Videotape.

Coughran:1983:CCA

[CGR83] W. M. Coughran, Jr., E. H.

Grosse, and D. J. Rose. CAzM: A circuit-analyzer with macromodeling. IEEE Transactions on Electron Devices, ED-30:1207– 1213, 1983.

Coughran:1986:VDS

[CGR86] William M. Coughran, Jr., Eric Grosse, and Donald J. Rose. Variation diminishing splines in simulation. SIAM Journal on Scientific and Statistical Computing, 7 (2):696–705, April 1986. CODEN SIJCD4. ISSN 0196-5204.

Coughran:1987:ACCA

[CGR87] W. M. Coughran, Jr., Eric H.

Grosse, and Donald J. Rose. Aspects of computational circuit analysis. In W. Fichtner and M. Morf, editors, VLSI CAD Tools and Applications, pages 105–127 (of x + 552). Kluwer Academic Publishers Group, Dordrecht, The Netherlands, 1987. ISBN 0-89838-193-2. LCCN TK7874 .V5572 1987.

Coughran:1981:CIS

[CH81] W. M. Coughran, Jr. and J. G.

Herriot. An algorithm to construct complete interpolating splines. Computing Science Technical Report 101, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1981.

Coughran:1990:MAT

[CJ90] W. M. Coughran, Jr. and J. W.

Jerome. Modular algorithms for transient semiconductor device simulation, part I: Analysis of the outer iteration. In R. E. Bank, editor, Computational Aspects of VLSI Design with an Emphasis on Semiconductor Device Simulation, pages 107–149. American Mathematical Society, Providence, RI, USA, 1990.

Coughran:1977:NAPL

[Cou77] W. M. Coughran, Jr. A note concerning the construction of a numerical analysis program library. Technical Memo 107, SLAC Computing Services, 1977.

Coughran:1980:IBVP

[Cou80] W. M. Coughran, Jr. On the Approximate Solution of Hyperbolic Initial-Boundary Value Problems. PhD thesis, Stanford University Computer Science Department, 1980. STAN-CS-80-806.

Coughran:1984:NBCIBVP

[Cou84]  W. M. Coughran, Jr.  On

noncharacteristic boundary conditions for discrete hyperbolic initial-boundary-value problems. Computing Science Technical Report 94, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1984.

Coughran:1985:NBC

[Cou85]  W. M. Coughran, Jr.  On

noncharacteristic boundary conditions for discrete hyperbolic initial-boundary-value problems. Journal of Computational Physics, 60(1):135–154, August 1985. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/002199918590021X.

Coughran:1997:NBS

[Cou97] W. M. Coughran, Jr. Networkbased scientific computation via Inferno. In Boisvert [Boi97], pages 267–269. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

Coughran:1988:CMSSCLS

[CPS88] W. M. Coughran, Jr., M. R.

Pinto, and R. K. Smith. Computational methods for steady-state CMOS latchup simulation. IEEE Transactions on Computer-Aided Design, 7:307–323, 1988.

Coughran:1989:CMS

[CPS89] W. M. Coughran, Jr., M. R.

Pinto, and R. K. Smith. Continuation methods in semiconductor device simulation. Journal of Computational and Applied Mathematics, 26(1–2):47– 65, June 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.math.utah.
edu/pub/bibnet/authors/c/coughran-
william-m.bib;   http://www.
sciencedirect.com/science/
article/pii/0377042789901477.

Coughran:1991:AGG

[CPS91] W. M. Coughran, Jr., M. R.

Pinto, and R. K. Smith. Adaptive grid generation for VLSI device simulation. IEEE Transactions on Computer-Aided Design, 10:1259–1275, 1991.

Coughran:1973:EOM

[CRSM73] W. M. Coughran, Jr., J. Rose, T. Shibuya, and V. McKoy. Equations-of-motion method: Potential energy curves for N_{2}, CO, C_{2}H_{4}. Journal of Chemical Physics, 58:2699–2709, 1973. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Coughran:1990:FDM

[CS90] W. M. Coughran, Jr. and N. L. Schryer. Faster device modeling using adaptive spatial meshes and continuation. In Workshop on Numerical Modeling of Processes and Devices for Integrated Circuits: NUPAD III, Technical Digest, pages 85–86. IEEE Circuits and Systems Society and Electron Devices Society, IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 1990. Extended abstract.

Grosse:1983:PPL

[GC83] E. H. Grosse and W. M. Coughran, Jr. The pine programming language. Numerical Analysis Manuscript 83-4, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1983. URL

ftp://netlib.bell-labs.com/
netlib/att/cs/doc/92/pine.
ps.gz.

Greenside:1984:NPF

[GC84] H. S. Greenside and W. M. Coughran, Jr. Nonlinear pattern formation near the onset of Rayleigh-Bénard convection. Physical Review A, 30:398– 428, 1984. CODEN PLRAAN.

ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Greenside:1988:MFOa

[GCC88] H. S. Greenside, M. C. Cross, and W. M. Coughran, Jr. Mean flows and the onset of chaos in largecell convection. Physical Review Letters, 60:2269–2272, 1988. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Greenside:1982:NPF

[GCS82] H. S. Greenside, W. M. Coughran, Jr., and N. L. Schryer. Nonlinear pattern formation near the onset of Rayleigh-Bénard convection. Physical Review Letters, 49:726–729, 1982. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Lamb:1990:OOA

[LFC90] P. Lamb, W. Fichtner, and W. M. Coughran, Jr. An objectoriented approach to mixedmode electrical circuit simulation. In A. Guasch, editor, Proceedings of the SCS Multiconference on Object-Oriented Simulation, pages 43–48. Society for Computer Simulation, 1990.

Pinto:1992:TDCa

[PBR^{+}92] M. R. Pinto, D. M. Boulin, C. S.

Rafferty, R. K. Smith, W. M. Coughran, Jr., I. C. Kizilyalli, and M. J. Thoma. Threedimensional characterization of bipolar transistors in a submicron BiCMOS technology using integrated process and device simulation. In IEDM Technical Digest ’92, pages 923–926. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992.

Pinto:1991:DSSa

[PCR^{+}91] M. R. Pinto, W. M. Coughran, Jr., C. S. Rafferty, E. Sangiorgi, and R. K. Smith. Device simulation for silicon ULSI. In K. Hess, J. P. Leburton, and U. Ravaioli, editors, Computational Electronics: Semiconductor Transport and Device Simulation, pages 3–13. Kluwer Academic Publishers, 1991.

Smith:1994:CCSa

[SC94] R. K. Smith and W. M. Coughran, Jr. Computational challenges in simulations of ULSI semiconductor devices. In T. N. Mudge and B. D. Shriver, editors, Proceedings of the 27th Hawaii International Conference on System Sciences, pages 7–15. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

Smith:1986:SASa

[SCF^{+}86] R. K. Smith, W. M. Coughran, Jr., W. Fichtner, D. J. Rose, and R. E. Bank. Some aspects of semiconductor device simulation. In R. Glowinski and J.-L. Lions, editors, Computing Methods in Applied Sciences and Engineering, VII, pages 3–12. North Holland-Elsevier Science Publishers, Amsterdam, New York, Oxford, 1986.